i am eight times older than my daughter now. in three years, i will be five times older than she is. how old am i?
Answer by DrBeeee(684)

Let m = my present age
d = daughter's present age
Presently
(1) m = 8*d
In 3 years I will be m + 3 and daughter d + 3 and at that time I will be 5 times as old as my daughter or
m + 3 = 5*(d + 3) or
m = 5d + 15 - 3
(2) m = 5d +12
Using (1) in (2) we have
8d = 5d + 12
3d = 12
d = 4
m = 32
In 3 years I will be 35 and my daughter will be 7. I will be 5 times as old as stated.
Answer is: I am 32 and my daughter is 4.